A state highway department is considering a bypass loop that is expected to save motorists $815,000 per year in gasoline and other automobile-related expenses. However, local businesses will experience revenue losses estimated to be $149,000 each year. The cost of the loop will be $8,600,000. Calculate the conventional B/C ratio using an interest rate of 8% per year and a 25-year project period.
Answer #1

Present value of net benefits = (815000 - 149000) * (P/A, 8%,25)

= 666000 * 10.674776

= 7109400.94

B/C ratio = Present value of net benefits / Initial cost

= 7109400.94 / 8600000

= 0.83

As B/C ratio is less than 1, this project should not be selected